# 1in1in1in1in

### v2

@@ Abstract
In this paper, we study the group of homeomorphisms of a surface that preserves the set of continuously differentiable curves. We give necessary and sufficient local conditions for a homeomorphism to be an element of this group. We also provide examples of several specific types of elements that are not diffeomorphisms.
@@ 1 Introduction
Given a smooth surface $S$, we say that a simple closed curve $\alpha$ in $S$ is $C^{1}$ if it is a properly embedded one-dimensional $C^{1}$ submanifold of $S$. Equivalently, these can also be defined as the image of a continuously differentiable and injective $\gamma:S^{1}\rightarrow S$ with $\gamma^{\prime}(t)\neq 0$ for all $t\in S^{1}$. We define the group $\operatorname{Homeo^{1}}(S)$ to be the set of $f\in\operatorname{Homeo}(S)$ such that both $f$ and $f^{-1}$ map every $C^{1}$ curve to a $C^{1}$ curve. In a separate paper [Boo24], we show that when $S$ is a compact orientable surface with genus at least 2, $\operatorname{Homeo^{1}}(S)$ is naturally isomorphic to the automorphisms of the $C^{1}$-curve graph.
The group of $C^{1}$ diffeomorphisms of $S$, denoted $\operatorname{Diff}^{1}(S)$, is a subgroup of $\operatorname{Homeo^{1}}(S)$. We show in Section 2 that this containment is proper by adapting an example of Le Roux–Wolff [LW22]. Since there are homeomorphisms that do not map every $C^{1}$ curve to a $C^{1}$ curve, we have the following inclusions:
@@ 2. $(ii)$
    the pushforward map $df_{p_{2}}:T_{p_{2}}S\rightarrow T_{f(p_{2})}S$ is well-defined and maps every element in $T_{p_{2}}S$ to 00,
@@ 1.2 The projective tangent bundle and transverse sequences
Figure 1: A transverse sequence of tangent lines from $y=x^{2}\sin(1/x)$ with slope 1 that converge to the origin along the $x$-axis.
@@ 1.3 Further questions
The previous two questions have been studied for both $\operatorname{Homeo}(S)$ and $\operatorname{Diff}(S)$. For compact $S$, they have both been answered positively, but the tools used for $\operatorname{Homeo}(S)$ are different from those used for $\operatorname{Diff}(S)$. Further discussion about these properties and others can be found in an article by Mann [Man21].
@@ 1.4 Paper structure and overview of the proof
Our paper is split into three parts. The first part is Section 2, where we give several types of examples showing Propositions 1.1 and 1.2 stated above. We then prove our Main Theorem. The proof is split into two parts, each corresponding to one of the logical directions of the statement. The forward direction involves recovering each of the three properties stated in the theorem from an element of $\operatorname{Homeo^{1}}(S)$, while the reverse direction requires us to show that any map with these properties is an element of $\operatorname{Homeo^{1}}(S)$.
@@ 1.5 Acknowledgments
[…]
In this section, we prove Propositions 1.1 and 1.2 from Section 1. We accomplish this by giving several types of maps that are elements of $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right){\setminus}\operatorname{Diff}^{1}\!\left(\mathbb{R}^{2}\right)$. The primary goal for many of these examples is to dismiss potentially simpler characterizations for elements of $\operatorname{Homeo^{1}}(S)$. Since the properties of $\operatorname{Homeo^{1}}(S)$ from our Main Theorem are dependent primarily on the local conditions around a point, we give these examples explicitly on $\mathbb{R}^{2}$.
@@ 2.1 Descriptions of the examples
In this section, we start with the examples for parts $(i)$ and $(ii)$ from Proposition 1.1. We then give a non-example that proves Proposition 1.2. Finally, we conclude this section with the examples for parts $(iii)$ and $(iv)$ from Propostion 1.1.
@@
This example proves the following proposition related to part $(i)$ from Proposition 1.1.
Proposition 2.1.
@@
We build $G:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ similar to an example given by Le Roux–Wolff [LW22, Section 5]. In their example, they change $y$-coordinates based on the slope of the line to the origin. We modify their construction by modifying both coordinates instead of only one of them.
Let $g:\mathbb{R}\rightarrow\mathbb{R}^{+}$ be a smooth homeomorphism such that $g(1)\neq 1$ and $g(t)=1$ whenever $t$ is outside of the interval $[1/2,2]$. Define the map $G:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ by
@@
Note that $G$ is a $C^{1}$ diffeomorphism at every point except the origin. We can show by direct computation that every $C^{1}$ curve through the origin is mapped by $G$ to a $C^{1}$ curve. Since $G^{-1}$ is of the same form as $G$, then $G^{-1}$ also maps every $C^{1}$ curve to a $C^{1}$ curve. Thus $G\in\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right)$.
$\xrightarrow{dG_{(0,0)}}$      Refer to caption
Figure 2: The action of $G$ on the vectors $(1,0)$, $(0,1)$, and $(1,1)$
@@
For part $(ii)$ from Proposition 1.1, this example proves the following:
Proposition 2.2.
There exists an element $H\in\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right){\setminus}\operatorname{Diff}^{1}\!\left(\mathbb{R}^{2}\right)$ fixing the origin such that the pushforward map $dH_{(0,0)}:T_{(0,0)}\mathbb{R}^{2}\rightarrow T_{(0,0)}\mathbb{R}^{2}$ is well-defined and maps every element in $T_{(0,0)}\mathbb{R}^{2}$ to 00.
@@
For any unit-speed $C^{1}$ parameterization of a curve $\gamma(t)$ such that $\gamma(0)=(0,0)$, the curve $H\circ\gamma$ has a tangent vector with magnitude 00 at the origin. By reparameterizing by $s=sgn(t)\,t^{2}$, it can be shown that the image of $H\circ\gamma$ is still a $C^{1}$ curve.
@@
Proving Proposition 1.2 : A cautionary non-example for $\mathbf{\operatorname{\textbf{Homeo}^{\textbf{1}}}(S)}$
@@
But $W^{-1}$ does not map every $C^{1}$ curve to a $C^{1}$ curve. At the origin, $\bar{d}W_{(0,0)}$ compresses the angles near 00 and $\pi$ in such a way that $W$ sends the curve defined by $y=x^{2}\sin(1/x)$ to a $C^{1}$ curve. Thus $W$ is not an element of $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right)$. ∎
@@
We return back to examples of $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right)$. Recall that elements of $\operatorname{Homeo^{1}}(S)$ induce a map on $\mathbb{P}TS$. By property $(b)$ from our Main Theorem, we know that this map on $\mathbb{P}TS$ must be a bijection. But the following result related to part $(iii)$ of Proposition 1.1 shows that this induced map does not have to be continuous.
Proposition 2.3.
@@
A natural question to consider is whether elements of $\operatorname{Homeo^{1}}(S)$ can have infinitely many non-differentiable points that accumulate on the surface. The following result related to part $(iv)$ from Proposition 1.1 gives an answer to this question.
Proposition 2.4.
@@
Consider an element $P_{0}$ from $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right){\setminus}\operatorname{Diff}^{1}\!\left(\mathbb{R}^{2}\right)$ similar to examples $(i)$ or $(iii)$ that is the identity on lines through the origin with slopes between $[-1,1]$ but not the identity along the $y$-axis. Let $\rho:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ be a smooth function away from $(0,0)$ such that $\rho(x,y)=P_{0}(x,y)$ when $||(x,y)||<\frac{1}{2}$ and $\rho(x,y)=(x,y)$ when $||(x,y)||\geq 1$. Let
$\rho_{n}(x,y)=2^{-(n+2)}\rho\left(2^{n+2}(x-2^{-n},y)\right)+(2^{-n},0)$
Note that each $\rho_{n}$ has a single non-differentiable point and support contained in the open ball of radius $2^{-(n+2)}$ around the point $x_{n}=(2^{-n},0)$. Note that these supports are all disjoint. Define $P:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ by
$P(x,y)=\left\{\begin{array}[]{ll}\rho_{n}(x,y)&\mbox{when }(x,y)\in\operatorname{supp}(\varphi_{n})\\ (x,y)&\mbox{otherwise}\end{array}\right.$
Any curve that is $C^{1}$ and passes through only finitely many of the $x_{n}$ will continue to be $C^{1}$ after $P$ is applied since each of the non-differentiable points are isolated. If a $C^{1}$ curve passes through infinitely many of the $x_{n}$, then it must have a slope of 0 at the origin. Moreover, by continuity of tangent lines, the slopes of this curve at the $x_{n}$ must be converging to 0 as $n\rightarrow\infty$. So $P$ will be the identity on this curve except at finitely many of the $x_{n}$. Thus $P$ maps every $C^{1}$ curve to a $C^{1}$ curve.
@@
In addition, this construction works whether or not the point $x_{\infty}=(0,0)$ supported on all of $\mathbb{R}^{2}$ is included. So the set of non-differentiable points for elements of $\operatorname{Homeo^{1}}(S)$ does not have to be a closed set nor finite.
2.2 The proof of Proposition 1.1
@@
$f(p)=\left\{\begin{array}[]{ll}(\varphi_{1}^{-1}\circ G\circ\varphi_{1})(p)&\mbox{ if }p\in U_{1}\\ (\varphi_{2}^{-1}\circ H\circ\varphi_{2})(p)&\mbox{ if }p\in U_{2}\\ (\varphi_{3}^{-1}\circ Q\circ\varphi_{3})(p)&\mbox{ if }p\in U_{3}\\ (\varphi_{4}^{-1}\circ P\circ\varphi_{4})(p)&\mbox{ if }p\in U_{4}\\ p&\mbox{ otherwise}\end{array}\right.$
By Propositions 2.1-2.4, $f\in\operatorname{Homeo^{1}}(S){\setminus}\operatorname{Diff}^{1}(S)$ and has the desired properties. ∎
@@ 3.1 A well-defined map of projective tangent space
Refer to caption $\alpha$$\beta$$\gamma$$\delta$$p$
Figure 8: Two topologically transverse $C^{1}$ curves $\alpha$ and $\beta$ at the point $p$ with the one-sided tangent curves $\gamma$ and $\delta$.
@@
By continuity, the origin must be a point on $\alpha$. Let $\alpha(t):\mathbb{R}\rightarrow\mathbb{R}^{2}$ be a $C^{1}$ parameterization of $\alpha$ such that $\alpha(0)=(0,0)$ and $\alpha^{\prime}(t)\neq 0$. Up to replacing $t$ by $-t$, we can assume that infinitely many of the $x_{n}$ occur at positive $t$ values. Let $x_{n_{k}}$ be the subsequence of of the $x_{n}$ that have positive $t$ values on $\alpha(t)$. Similarly, there is a $C^{1}$ parameterization $\beta(s):\mathbb{R}\rightarrow\mathbb{R}^{2}$ of $\beta$ with $\beta(0)=0$ and $\beta^{\prime}(s)\neq 0$. Let $t_{k}$ and $s_{k}$ be sequences such that $\alpha(t_{k})=x_{n_{k}}=\beta(s_{k})$.
@@
Thus $\alpha^{\prime}(0)$ and $\beta^{\prime}(0)$ must lie on the same line, which is the direction that the $x_{n_{k}}$ approach 00. Since both $\alpha^{\prime}(0)$ and $\beta^{\prime}(0)$ are nonzero, then

### v3

# Homeomorphisms of surfaces that preserve continuously differentiable curves

@@ Abstract
In this paper, we study $\operatorname{Homeo}^{1}(S)$, the group of homeomorphisms of a surface that preserve the set of one-dimensional $C^{1}$ submanifolds of that surface. The group $\operatorname{Homeo}^{1}(S)$ belongs to a family of similarly defined groups $\operatorname{Homeo}^{k}(S)$ that were recently introduced by the author. In a separate paper, we have shown that for most closed surfaces, $\operatorname{Homeo}^{k}(S)$ is naturally isomorphic to the automorphisms of a smooth fine curve graph. By contrast, the work in this paper gives local conditions that characterize $\operatorname{Homeo}^{1}(S)$. We show that there exists a collection of conditions that are both necessary and sufficient for a homeomorphism of the surface to be an element of this group. These conditions primarily depend upon the structure of the induced map on the projective tangent bundle. Additionally, we provide examples of several types of elements of $\operatorname{Homeo}^{1}(S)$ that are not diffeomorphisms. These include inducing discontinuous maps on the projective tangent bundle and having infinitely many non-differentiable points.
pacs:
MSC Classification]57K20, 20F65, 57S05
@@ 1 Introduction
Given a smooth surface $S$, we say that a simple closed curve $\alpha$ in $S$ is $C^{1}$ if it is a properly embedded one-dimensional $C^{1}$ submanifold of $S$. Equivalently, these can also be defined as the image of a continuously differentiable and injective $\gamma:S^{1}\rightarrow S$ with $\gamma^{\prime}(t)\neq 0$ for all $t\in S^{1}$. We define the group $\operatorname{Homeo^{1}}(S)$ to be the set of $f\in\operatorname{Homeo}(S)$ such that both $f$ and $f^{-1}$ map every $C^{1}$ curve to a $C^{1}$ curve. In a separate paper [1], we show that when $S$ is a compact orientable surface with genus at least 2, $\operatorname{Homeo^{1}}(S)$ is naturally isomorphic to the automorphisms of the $C^{1}$-curve graph.
The group of $C^{1}$ diffeomorphisms of $S$, denoted $\operatorname{Diff}^{1}(S)$, is a subgroup of $\operatorname{Homeo^{1}}(S)$. We show in Section 2 that this containment is proper by adapting an example of Le Roux–Wolff [2]. Since there are homeomorphisms that do not map every $C^{1}$ curve to a $C^{1}$ curve, we have the following inclusions:
@@ 2. $(ii)$
    the pushforward map $df_{p_{2}}:T_{p_{2}}S\rightarrow T_{f(p_{2})}S$ is well-defined and maps every element in $T_{p_{2}}S$ to 0,
@@ 1.2 The projective tangent bundle and transverse sequences
Figure 1: A transverse sequence of tangent lines from $y=x^{2}\sin(1/x)$ with slope 1 that converge to the origin along the $x$-axis
@@ 1.3 Further questions
The previous two questions have been studied for both $\operatorname{Homeo}(S)$ and $\operatorname{Diff}(S)$. For compact $S$, they have both been answered positively, but the tools used for $\operatorname{Homeo}(S)$ are different from those used for $\operatorname{Diff}(S)$. Further discussion about these properties and others can be found in an article by Mann [3].
@@ 1.4 Paper structure and overview of the proof
Our paper is split into three parts. The first part is Section 2, where we give several types of examples showing Propositions 2 and 3 stated above. We then prove our Main Theorem. The proof is split into two parts, each corresponding to one of the logical directions of the statement. The forward direction involves recovering each of the three properties stated in the theorem from an element of $\operatorname{Homeo^{1}}(S)$, while the reverse direction requires us to show that any map with these properties is an element of $\operatorname{Homeo^{1}}(S)$.
@@ 1.5 Acknowledgments
2 Example elements from $\mathbf{\operatorname{\textbf{Homeo}}^{1}(S)}$
In this section, we prove Propositions 2 and 3 from Section 1. We accomplish this by giving several types of maps that are elements of $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right){\setminus}\operatorname{Diff}^{1}\!\left(\mathbb{R}^{2}\right)$. The primary goal for many of these examples is to dismiss potentially simpler characterizations for elements of $\operatorname{Homeo^{1}}(S)$. Since the properties of $\operatorname{Homeo^{1}}(S)$ from our Main Theorem are dependent primarily on the local conditions around a point, we give these examples explicitly on $\mathbb{R}^{2}$.
@@ 2.1 Descriptions of the examples
In this section, we start with the examples for parts $(i)$ and $(ii)$ from Proposition 2. We then give a non-example that proves Proposition 3. Finally, we conclude this section with the examples for parts $(iii)$ and $(iv)$ from Propostion 2.
@@
This example proves the following proposition related to part $(i)$ from Proposition 2.
Proposition 4.
@@
We build $G:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ similar to an example given by Le Roux–Wolff [2, Section 5]. In their example, they change $y$-coordinates based on the slope of the line to the origin. We modify their construction by modifying both coordinates instead of only one of them.
Let $g:\mathbb{R}\rightarrow\mathbb{R}^{+}$ be a smooth map such that $g(1)\neq 1$ and $g(t)=1$ whenever $t$ is outside of the interval $[1/2,2]$. Define the map $G:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ by
@@
Note that $G$ is a $C^{1}$ diffeomorphism at every point except the origin. We can show by direct computation that every $C^{1}$ curve through the origin is mapped by $G$ to a $C^{1}$ curve. Since $G^{-1}$ is of the same form as $G$ by replacing $g$ by $1/g$, then $G^{-1}$ also maps every $C^{1}$ curve to a $C^{1}$ curve. Thus $G\in\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right)$.
Figure 2: The action of $G$ on the vectors $(1,0)$, $(0,1)$, and $(1,1)$ based at the origin
@@
For part $(ii)$ from Proposition 2, this example proves the following:
Proposition 5.
There exists an element $H\in\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right){\setminus}\operatorname{Diff}^{1}\!\left(\mathbb{R}^{2}\right)$ fixing the origin such that the pushforward map $dH_{(0,0)}:T_{(0,0)}\mathbb{R}^{2}\rightarrow T_{(0,0)}\mathbb{R}^{2}$ is well-defined and maps every element in $T_{(0,0)}\mathbb{R}^{2}$ to 0.
@@
For any unit-speed $C^{1}$ parameterization of a curve $\gamma(t)$ such that $\gamma(0)=(0,0)$, the curve $H\circ\gamma$ has a tangent vector with magnitude 0 at the origin. By reparameterizing by $s=sgn(t)\,t^{2}$, it can be shown that the image of $H\circ\gamma$ is still a $C^{1}$ curve.
@@
Proving Proposition 3 : A cautionary non-example for $\mathbf{\operatorname{\textbf{Homeo}^{\textbf{1}}}(S)}$
@@
But $W^{-1}$ does not map every $C^{1}$ curve to a $C^{1}$ curve. At the origin, $\bar{d}W_{(0,0)}$ compresses the angles near 0 and $\pi$ in such a way that $W$ sends the curve defined by $y=x^{2}\sin(1/x)$ to a $C^{1}$ curve. Thus $W$ is not an element of $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right)$. ∎
@@
We return back to examples of $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right)$. Recall that elements of $\operatorname{Homeo^{1}}(S)$ induce a map on $\mathbb{P}TS$. By property $(b)$ from our Main Theorem, we know that this map on $\mathbb{P}TS$ must be a bijection. But the following result related to part $(iii)$ of Proposition 2 shows that this induced map does not have to be continuous.
Proposition 6.
@@
A natural question to consider is whether elements of $\operatorname{Homeo^{1}}(S)$ can have infinitely many non-differentiable points that accumulate on the surface. The following result related to part $(iv)$ from Proposition 2 gives an answer to this question.
Proposition 7.
@@
We start with Le Roux–Wolfe’s [2] example element from $\operatorname{Homeo^{1}}\!\left(\mathbb{R}^{2}\right){\setminus}\operatorname{Diff}^{1}\!\left(\mathbb{R}^{2}\right)$:
$P_{0}(x,y)=\left\{\begin{array}[]{ll}(x,x\,p_{0}(y/x))&\mbox{if }x\neq 0\\ (x,y)&\mbox{if }x=0\end{array}\right$.
where $p_{0}:\mathbb{R}\rightarrow\mathbb{R}$ is a diffeomorphism such that $p_{0}(x)=x$ when $x$ is outside $[1/2,2]$, but $p_{0}(1)\neq 1$. Our goal is to make a function that has similar behavior at a sequence of points that converge to the origin, while becoming increasingly nicer as the points make the approach. To that end, we define a sequence of functions
$p_{k}(x)=\frac{1}{2^{k}}p_{0}(x)+\left(1-\frac{1}{2^{k}}\right)x$
Note that this family approaches the identity as $k\rightarrow\infty.$ Let $\varphi_{0}:\mathbb{R}^{2}\rightarrow[0,1]$ be a smooth bump function such that $\varphi_{0}(x,y)=1$ when $(x,y)||<\frac{1}{2}$ and $\varphi_{0}(x,y)=0$ when $(x,y)||\geq 1$. Similarly, define a sequence of functions
$\varphi_{k}(x,y)=\varphi_{0}\left(2^{k+2}(x-2^{-k}),2^{k+2}y\right)$
We can now define the functions
$\rho_{n}(x,y)=\left(x,\left(x-2^{-n}\right)p_{n}\left(\frac{y}{x-2^{-n}}\right)\varphi_{n}(x,y)+\left(1-\varphi_{n}(x,y)\right)y\right)$
@@
Note that each $\rho_{n}$ has a single non-differentiable point and support contained in the open ball of radius $2^{-(n+2)}$ around the point $x_{n}=(2^{-n},0)$, which means that these supports are all disjoint. Define $P:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ by
$P(x,y)=\left\{\begin{array}[]{ll}\rho_{n}(x,y)&\mbox{when }(x,y)\in\operatorname{supp}(\rho_{n})\\ (x,y)&\mbox{otherwise}\end{array}\right$.\\
Any curve that is $C^{1}$ and passes through only finitely many of the $\operatorname{supp}(\rho_{n})$ will continue to be $C^{1}$ after $P$ is applied since each of the non-differentiable points are isolated. On the other hand, it can be shown by direct computation that the image of any $C^{1}$ curve that passes through infinitely many of the $\operatorname{supp}(\rho_{n})$ will still be a $C^{1}$ curve. In particular, when the $p_{n}$ approach the identity, the distortion on the curve also decreases as it approaches the origin, preserving the continuity of the tangent directions. Thus $P$ maps every $C^{1}$ curve to a $C^{1}$ curve.
@@
In addition, this construction also works when point $x_{\infty}=(0,0)$ is non-differentiable and has support on the region of $\mathbb{R}^{2}$ where $|y|>|x|$. So the set of non-differentiable points for elements of $\operatorname{Homeo^{1}}(S)$ does not have to be a closed set nor finite.
2.2 The proof of Proposition 2
@@
[…]
By Propositions 4-7, $f\in\operatorname{Homeo^{1}}(S){\setminus}\operatorname{Diff}^{1}(S)$ and has the desired properties. ∎
@@ 3.1 A well-defined map of projective tangent space
Figure 8: Two topologically transverse $C^{1}$ curves $\alpha$ and $\beta$ at the point $p$ with the one-sided tangent curves $\gamma$ and $\delta$
@@
By continuity, the origin must be a point on $\alpha$. Let $\alpha(t):\mathbb{R}\rightarrow\mathbb{R}^{2}$ be a $C^{1}$ parameterization of $\alpha$ such that $\alpha(0)=(0,0)$ and $\alpha^{\prime}(t)\neq 0$. Up to replacing $t$ by t-t, we can assume that infinitely many of the $x_{n}$ occur at positive $t$ values. Let $x_{n_{k}}$ be the subsequence of of the $x_{n}$ that have positive $t$ values on $\alpha(t)$. Similarly, there is a $C^{1}$ parameterization $\beta(s):\mathbb{R}\rightarrow\mathbb{R}^{2}$ of $\beta$ with $\beta(0)=0$ and $\beta^{\prime}(s)\neq 0$. Let $t_{k}$ and $s_{k}$ be sequences such that $\alpha(t_{k})=x_{n_{k}}=\beta(s_{k})$.
@@
Thus $\alpha^{\prime}(0)$ and $\beta^{\prime}(0)$ must lie on the same line, which is the direction that the $x_{n_{k}}$ approach 0. Since both $\alpha^{\prime}(0)$ and $\beta^{\prime}(0)$ are nonzero, then